### v1

@@ Abstract
Can the causal effects estimated in experiment be generalized to real-world scenarios? This question lies at the heart of social science studies. External validity primarily assesses whether experimental effects persist across different settings, implicitly presuming the experiment’s ecological validity—that is, the consistency of experimental effects with their real-life counterparts. However, we argue that this presumed consistency may not always hold, especially in experiments involving multi-dimensional decision processes, such as conjoint experiments. We introduce a formal model to elucidate how attention and salience effects lead to three types of inconsistencies between experimental findings and real-world phenomena: amplified effect magnitude, effect sign reversal, and effect importance reversal. We derive testable hypotheses from each theoretical outcome and test these hypotheses using data from various existing conjoint experiments and our own experiments. Drawing on our theoretical framework, we propose several recommendations for experimental design aimed at enhancing the generalizability of survey experiment findings.
@@ Introduction
The randomized experiment is often hailed as the gold standard for identifying and estimating causal effects within the social sciences. One popular such method is conjoint experiment, which is frequently employed by researchers to examine how individual preferences for candidates or immigrants are influenced by characteristics such as gender, race, experience, and skills (Eggers, Vivyan and Wagner, 2018; Carnes and Lupu, 2016; Sanbonmatsu, 2002). However, the interest of researchers extends well beyond the mere observation of causal effects within an experimental setting. Especially, we ask whether these experimental estimates accurately mirror real-world effects. For instance, in a conjoint survey experiment that reveals a preference for female candidates with an average causal effect of approximately 2%, researchers hope that this finding corresponds to an actual increase in preference for female candidates in real-world scenarios. When such a correspondence is established, we regard the experimental effect as being consistent with the real-world effect. This consistency, which establishes the ecological validity of experiment, serves as a crucial concern of social science studies.
Ecological validity represents a critical aspect of external validity, distinct from internal validity, which pertains to the validity of conclusions within the experimental setting (Kihlstrom 2021). As depicted in Figure 1, external validity primarily concerns the generalization of results from the experimental domain to real-world contexts with varied populations, treatments, outcomes, and contexts (Cook, Campbell and Shadish 2002). A fundamental premise is that the effects observed in an experiment should be consistent with real-life effects. If the effects estimated in the experiment do not align with those experienced by the same individuals in the real world, it undermines the rationale for discussing the generalizability of these results to other populations, treatments, outcomes, or contexts.
Figure 1: Experimental Validity: Ecological validity as a premise or a special subset of External Validity
We demonstrate that experiments involving multi-dimensional decision-making processes, such as the widely used vignette and conjoint survey experiments with multiple attributes manipulated at random, frequently fail to achieve ecological validity. This failure raises substantial concerns about our ability to generalize experimental insights to other real-life scenarios effectively. Within the scope of this paper, we specifically examine three types of consistency: effect magnitude, effect direction, and attribute importance. The Consistency of Effect Magnitude asserts that the observed size of an experimental effect should mirror the magnitude (and the sign) observed in real-world phenomena. This criterion, while offering the deepest insights into real-world applicability, stands as the most challenging to fulfill. The Consistency of Effect Direction posits that the sign (positive or negative) of an experimental effect should match its real-world counterpart. Achieving magnitude consistency ensures direction consistency by default, rendering the latter a less rigorous yet still crucial requirement for real-world insights. Lastly, Consistency of Attribute Importance emphasizes the need for experimental outcomes to truthfully represent the relative magnitude of different attributes. This aspect is particularly valuable for researchers aiming to identify which attributes most significantly influence variations in outcomes, marking it as another derivative of magnitude consistency.
@@
We develop a formal theory to elucidate why experiments involving multi-dimensional decision-making frequently fall short of meeting the three consistency criteria. At the heart of our theoretical framework lies the interplay between attention and salience. Conventional decision-making theories have often posited that individuals possess unlimited attention and exhibit consistent salience across contexts (i.e. they take account for all possible factors and the weight for each factor is independent or unchanged in different situations). While these assumptions yield straightforward theoretical predictions, they have been increasingly called into question due to numerous observed anomalies and paradoxes. Notably, the Allais (1953) and Ellsberg (1961) paradoxes serve as prominent examples that challenge these traditional views. The former one shows that individual frequently violate the independence axiom of expected utility theory and the latter one shows that people averse ambiguity and their decisions are inconsistent with subjective expected utility theory. While the bulk of existing research has focused on refining decision theory to account for these discrepancies, our work pioneers the application of a behavioral theoretical approach to systematically address these inconsistencies within experimental research methodologies.
@@
We derive several corresponding testable hypotheses from our theoretical results. For example, a key hypothesis is that the observed effect diminishes or even reverses as the number of attributes increases. To empirically test these hypotheses, we analyze data from multiple existing studies and conduct additional, pre-registered conjoint experiments  [^4] The study is pre-registered and approved by authors’ Institutional Review Board; More details are in the SI D.. The experimental findings consistently support our hypotheses, thereby reinforcing our theoretical framework. We do not claim that our proposed theory is the only reason that generates those inconsistency. However, from our theory, we can derive some specific hypotheses that may not straightforwardly derived from other theories. We believe that the positive results from testing those hypotheses could further support our theory. In light of mechanisms driving inconsistencies identified in our theory, we propose several experimental design recommendations in the discussion section aimed at alleviating these issues.
Our research intersects with and contributes to several vital streams of literature. At the core of our study is a pivotal inquiry into the conjoint experiment literature, a timely and critical exploration considering the burgeoning body of research on conjoint analysis since the seminal work of Hainmueller, Hopkins and Yamamoto (2014). Subsequent investigations have primarily honed in on statistical concerns—such as external validity (De la Cuesta, Egami and Imai, 2022), measurement errors (Clayton et al., 2023), power analysis (Stefanelli and Lukac, 2020), survey satisficing (Bansak et al., 2018, 2021), social desirability bias (Horiuchi, Markovich and Yamamoto, 2022), hypotheses testing (Ham, Imai and Janson, 2022), and multiple testing (Horiuchi, Markovich and Yamamoto, 2022)—yet, the fundamental issue of real-world consistency remains underexplored. Hainmueller, Hangartner and Yamamoto (2014) empirically find that survey experiments have high ecological validity [^5] In their paper, they primarily describe it as external validity. Actually, they demonstrate both external validity (comparing different individuals) and ecological validity (comparing experimental effects and real-world effects).. One main reason is that their experiment successfully eliminates the attention effect so that the consideration set of the experiment is almost the same to the real-world one. Our findings provide a theoretical explanation to understand conditions under which conjoint experiments are reliable.
@@
Third, we contribute to the discussion on external validity, which is a core concern in experimental research, especially in lab and survey experiments (Barabas and Jerit 2010; List and Levitt 2005; Gaines, Kuklinski and Quirk 2007). Slough and Tyson (2023, 2024) establish a formal framework to clarify the external validity. They identify conditions when different studies are target-equivalence and target-congruence. They particularly emphasize whether studies can be harmonized along treatment, control, and the measurement of the outcome. Egami and Hartman (2023) also develop a unified framework to discuss four types of external validity, including population, contexts, treatments, and outcomes, as proposed by Cook, Campbell and Shadish (2002). Several estimators are then proposed to conduct effect generalization and sign generalization. Most studies primarily focus on whether the results can be generalized to a new population (Mullinix et al. 2015; Huang 2022). Our paper focuses on ecological validity, which compares the experimental effect and real-world effect even for the same individuals, treatments, and outcomes. Many factors can affect ecological validity, such as the famous Hawthorne effect (Adair, 1984). We provide a new explanation to analyze ecological validity for experiments involving multi-dimensional decision processes.
@@ 1 Theoretical Model of Experiment
In practice, most conjoint experiments measure the forced choice outcome $Y$; that is, whether a candidate is chosen or not, denoted by $1$ or 00. Theoretically, the DM chooses candidate $A_{1}$ over $A_{2}$ because the DM has a higher evaluation for the former. Therefore, $Y$ is a function of the utility difference. The corresponding utility differences in World 1 are $V^{1}_{i}=V_{i}(A_{1})-V_{i}(A_{2})$, and in World 2, they are $V^{2}_{i}=V_{i}(\tilde{A}_{1})-V_{i}(A_{2})$  [^6] Note that $V_{i}(A_{2})$ in the two hypothetical worlds may differ due to the salience effect, which we will discuss in Section 3.. It is evident that the binary choice potential outcome can be defined as $Y_{i}^{j}=\mathbbm{1}[V^{j}_{i}\geq 0]$, where $j=1,2$.
Because the utility difference drives the potential outcome we observe, without loss of generality, throughout the main text, we focus on $V$ rather than $Y$  [^7] For outcomes that are preference scores, our results also hold because the score is also a function of the evaluation. In some survey experiments, the DM may only observe one alternative. In such cases, we can simply set $V_{i}(A_{2})=0$.. Consequently, the individual causal effect for attribute $X_{1}$ when it changes from $x_{1}$ to $\tilde{x}_{1}$, given the other attributes remain constant, is defined as the difference-in-differences $V^{1}_{i}-V^{2}_{i}$.
Definition 1 (ITE).
Given an alternative set $A=\{A_{1},\tilde{A_{1}},A_{2}\}$ characterized by attributes $X=(X_{1},X_{2},...,X_{K})$. Let $X^{A_{1}}=(x_{j},x_{-j})$ represent the attributes of $A_{1}$, where $x_{j}$ is the value of attribute $X_{j}$ for $A_{1}$ and $x_{-j}$ denotes the values of all other attributes except $X_{j}$. Similarly, let $\tilde{X}^{\tilde{A}_{1}}=(\tilde{x}_{j},x_{-j})$ and $X^{A_{2}}=(x^{\prime}_{j},x^{\prime}_{-j})$ represent the attributes of $\tilde{A}_{1}$ and $A_{2}$. For a DM $i$, the individual treatment effect of attribute $X_{j}$ with respect to changing its value from $x_{j}$ to $\tilde{x}_{j}$ is defined as: $ITE^{i}=[V_{i}(X^{A_{1}})-V_{i}(X^{A_{2}})]-[V_{i}(\tilde{X}^{A_{1}})-V_{i}(X^{A_{2}})]$.
This concept is also referred to as the individual component effect (ICE) by Abramson et al. (2023). The formulation aligns with methodologies employed in the conjoint experiment literature, as discussed by Hainmueller, Hopkins and Yamamoto (2014). Given that alternatives are characterized by attributes $X$, we also use the notation $ITE^{i}=[V_{i}(A_{1})-V_{i}(A_{2})]-[V_{i}(\tilde{A}_{1})-V_{i}(A_{2})]$ to denote ITE in this paper. This example of ITE underscores the notion that, within a multi-dimensional decision-making environment, the ITE is contingent not only on the treatment attribute $X_{1}$ but also on other attributes.
It is worth noting that the formal derivation of the ITE hinges on other assumptions: Stability and No Carryover Effects, and No Profile-Order Effects, which are standard in the conjoint experiment literature. A more formal discussion is available in the Supplementary Information (SI) 4. The Marginal Component Effect (MCE) is defined as a weighted sum of ITEs, and the Average Marginal Component Effect (AMCE) represents the expected value of MCE across the distribution of DMs.
@@ 2 Attention Effect
Individuals have limited attention spans. Economists have observed that when people make decisions, such as purchasing a computer, they do not consider all possible options (Hausman, 2008). Within the literature, the subset of alternatives that a DM considers when making a decision is referred to as the “consideration set” (Hausman, 2008; Wright and Barbour, 1977). It has been well-argued and demonstrated that due to cognitive limitations, humans cannot allocate sufficient attention to every potential attribute (Stigler, 1961; Jones and Baumgartner, 2005; Chetty, Looney and Kroft, 2009). This indicates that the consideration set is typically smaller than the full set $X$ that may influence a decision. We denote the consideration set in real-world decision-making as $X^{r}\subset X$ and the consideration set in an experimental context as $X^{e}\subset X$.
@@
The consideration set utilized in experiments, denoted as $X^{e}$, is often a subset of the real-world consideration set, $X^{r}$, for several reasons. Firstly, researchers seldom possess complete knowledge of which attributes constitute the real-world consideration set $X^{r}$. Secondly, even with such knowledge, it is impractical to present all relevant information about these attributes within the confines of a single survey experiment  [^8] A more general assumption could be that $X^{e}\neq X^{r}$. However, the empirical results in the paper provide little evidence on this assumption..
@@ 2.1 Inconsistency of Effect Magnitude
To compare the two causal effects without a salience effect, we establish a benchmark assumption that the salience ratio for any given dimension remains constant.
@@
This assumption reflects the natural idea that if, in one circumstance, a DM considers attribute $j$ (e.g., gender) to be more salient than attribute $k$ (e.g., age), that is, $\alpha_{j}>\alpha_{k}$, then in another similar circumstance, the same DM should still perceive gender as more salient than age, $\alpha^{\prime}_{j}>\alpha^{\prime}_{k}$, even if the exact values of salience could differ ($\alpha_{j}\neq\alpha^{\prime}_{j}$ and $\alpha_{k}\neq\alpha^{\prime}_{k}$). The assumption implies that their relative magnitude should remain the same. This assumption is not related to the consistency of relative importance because the latter is defined on the causal effect rather than salience.
@@
We will test this hypothesis in the following subsection. Specifically, following common practice, the experimental causal effect is measured by the AMCE. As noted in the SI 4, since AMCE is essentially a weighted sum of ITE, the amplification bias of ITE is also reflected in AMCE.
@@ 3 Salience Effect
Exploring further, we investigate the distortion in the linkage between experimental causal effects and real-world implications, moving beyond the assumption of fixed salience in the previous section. We adopt a psychological framework developed by Bordalo, Gennaioli and Shleifer (2012) to elucidate how variations in attribute salience can influence this relationship. This approach posits that individuals disproportionately weight more salient attributes when evaluating alternatives.

### v2

@@ Abstract
Can the causal effects estimated in an experiment be generalized to real-world scenarios? This question lies at the heart of social science studies. External validity primarily assesses whether experimental effects persist across different settings, implicitly presuming the consistency of experimental effects with their real-life counterparts. However, we argue that this presumed consistency may not always hold, especially in experiments involving multi-dimensional decision processes, such as conjoint experiments. We introduce a formal model to elucidate how attention and salience effects lead to three types of inconsistencies between experimental findings and real-world phenomena: amplified effect magnitude, effect sign reversal, and effect importance reversal. We derive testable hypotheses from each theoretical outcome and test these hypotheses using data from various existing conjoint experiments and our own experiments. Drawing on our theoretical framework, we propose several recommendations for experimental design aimed at enhancing the generalizability of survey experiment findings.
@@ Introduction
The randomized experiment is considered as the gold standard for identifying and estimating causal effects within the social sciences. One popular such method is the conjoint experiment, which is frequently employed by researchers to examine how individual preferences for candidates or immigrants are influenced by characteristics such as gender, race, experience, and skills (Eggers, Vivyan and Wagner, 2018; Carnes and Lupu, 2016; Sanbonmatsu, 2002). However, the interest of researchers extends well beyond the mere observation of causal effects within an experimental setting. Specifically, we ask whether these experimental estimates accurately mirror real-world effects. For instance, in a conjoint survey experiment that reveals a preference for female candidates with an average causal effect of approximately 2%, researchers hope that this finding corresponds to an actual increase in preference for female candidates in real-world scenarios. When such a correspondence is established, we regard the experimental effect as being consistent with the real-world effect.
We demonstrate that experiments involving multi-dimensional decision-making processes, such as the widely used vignette and conjoint survey experiments with multiple attributes manipulated at random, often fail to achieve such consistency. This failure raises substantial concerns about our ability to generalize experimental insights to other real-life scenarios effectively. Within the scope of this paper, we specifically examine three types of consistency: effect magnitude, effect direction, and attribute importance. The Consistency of Effect Magnitude asserts that the observed size of an experimental effect should mirror the magnitude (and the sign) observed in real-world phenomena. This criterion, while offering the deepest insights into real-world applicability, stands as the most challenging to fulfill. The Consistency of Effect Direction posits that the sign (positive or negative) of an experimental effect should match its real-world counterpart. Lastly, Consistency of Attribute Importance emphasizes the need for experimental outcomes to truthfully represent the relative magnitude of different attributes. This aspect is particularly valuable for researchers aiming to identify which attributes most significantly influence variations in outcomes.
@@
We develop a formal theory to elucidate why and under what conditions experiments involving multi-dimensional decision-making frequently fall short of meeting the three consistency criteria. At the heart of our theoretical framework lies the interplay between attention and salience. Conventional decision-making theories have often posited that individuals possess unlimited attention and exhibit consistent salience across contexts (i.e. they take account for all possible factors and the weight for each factor is independent or unchanged in different situations). While these assumptions yield straightforward theoretical predictions, they have been increasingly called into question due to numerous observed anomalies and paradoxes. Notably, the Allais (1953) and Ellsberg (1961) paradoxes serve as prominent examples that challenge these traditional views. The former one shows that individual frequently violate the independence axiom of expected utility theory and the latter one shows that people averse ambiguity and their decisions are inconsistent with subjective expected utility theory. While the bulk of existing research has focused on refining decision theory to account for these discrepancies, our work pioneers the application of a behavioral theoretical approach to systematically address these inconsistencies within experimental research methodologies.
@@
We derive several corresponding testable hypotheses from our theoretical results. For example, a key hypothesis is that the observed effect diminishes or even reverses as the number of attributes increases. To empirically test these hypotheses, we analyze data from multiple existing studies and conduct additional, pre-registered conjoint experiments. [^4] The study is pre-registered and approved by authors’ Institutional Review Board; More details are in the SI D. The experimental findings consistently support our hypotheses. We do not claim that our proposed theory is the only reason that generates those inconsistency. However, from our theory, we can derive some specific hypotheses that may not straightforwardly derived from other theories. We believe that the positive results from testing those hypotheses could further support our theory. In light of mechanisms driving inconsistencies identified in our theory, we propose several experimental design recommendations in the discussion section aimed at alleviating these issues.
Our research intersects with and contributes to several vital streams of literature. At the core of our study is a pivotal inquiry into the conjoint experiment literature, a timely and critical exploration considering the burgeoning body of research on conjoint analysis since the seminal work of Hainmueller, Hopkins and Yamamoto (2014). Subsequent investigations have primarily honed in on statistical concerns—such as external validity (De la Cuesta, Egami and Imai, 2022), measurement errors (Clayton et al., 2023), power analysis (Stefanelli and Lukac, 2020), survey satisficing (Bansak et al., 2018, 2021), social desirability bias (Horiuchi, Markovich and Yamamoto, 2022), hypotheses testing (Ham, Imai and Janson, 2022), and multiple testing (Horiuchi, Markovich and Yamamoto, 2022)—yet, the fundamental issue of real-world consistency remains underexplored. Hainmueller, Hangartner and Yamamoto (2014) empirically find that survey experiments have high ecological validity [^5] In their paper, they primarily describe it as external validity. Actually, they demonstrate both external validity (comparing different individuals) and ecological validity (comparing experimental effects and real-world effects).. One main reason is that their experiment successfully eliminates the attention effect so that the consideration set of the experiment is almost the same to the real-world one. Jenke et al. (2021) utilize eye-tracking to examine attention dynamics and the stability of conjoint experiments. Their findings offer empirical support for our assumptions and theoretical framework. Firstly, they observe that respondents’ attention diminishes markedly as the number of attributes increases. Secondly, while their research concludes that conjoint experiments are substantially stable, they still note some modest variations, including effect reversals. Our study provides a theoretical framework to understand the conditions under which conjoint experiments are reliable or unreliable.
@@
Third, we contribute to the discussion on external validity, which is a core concern in experimental research, especially in lab and survey experiments (Barabas and Jerit 2010; List and Levitt 2005; Gaines, Kuklinski and Quirk 2007). Slough and Tyson (2023, 2024) establish a formal framework to clarify the external validity. They identify conditions when different studies are target-equivalence and target-congruence. They particularly emphasize whether studies can be harmonized along treatment, control, and the measurement of the outcome. Egami and Hartman (2023) also develop a unified framework to discuss four types of external validity, including population, contexts, treatments, and outcomes, as proposed by Cook, Campbell and Shadish (2002). Several estimators are then proposed to conduct effect generalization and sign generalization. Most studies primarily focus on whether the results can be generalized to a new population (Mullinix et al. 2015; Huang 2022). Our paper focuses on the consistency between the experimental effect and real-world effect, even for the same individuals, treatments, and outcomes. Many factors can affect consistency, such as the famous Hawthorne effect (Adair, 1984). We provide a new explanation for experiments involving multi-dimensional decision processes.
@@ 1 Theoretical Model of Experiment
In practice, most conjoint experiments measure the forced choice outcome $Y$; that is, whether a candidate is chosen or not, denoted by $1$ or 00. Theoretically, the DM chooses candidate $A_{1}$ over $A_{2}$ because the DM has a higher evaluation for the former. Because the utility difference drives the potential outcome we observe, without loss of generality, throughout the main text, we focus on $V$ rather than $Y$  [^6] It is evident that the binary choice potential outcome can be defined as $Y_{i}^{j}=\mathbbm{1}[V^{j}_{i}\geq 0]$, where $j=1,2$. For outcomes that are preference scores, our results also hold because the score is also a function of the evaluation. In some survey experiments, the DM may only observe one alternative. In such cases, we can simply set $V_{i}(A_{2})=0$.. Therefore, $Y$ is a function of the utility difference. The corresponding utility differences in World 1 are $V^{1}_{i}=V_{i}(A_{1})-V_{i}(A_{2})$, and in World 2, they are $V^{2}_{i}=V_{i}(\tilde{A}_{1})-V_{i}(A_{2})$  [^7] Note that $V_{i}(A_{2})$ in the two hypothetical worlds may differ due to the salience effect, which we will discuss in Section 3.. Consequently, the individual causal effect for attribute $X_{1}$ when it changes from $x_{1}$ to $\tilde{x}_{1}$, given the other attributes remain constant, is defined as the difference-in-differences $V^{1}_{i}-V^{2}_{i}$.
This concept is also referred to as the individual component effect (ICE) by Abramson et al. (2023). The formulation aligns with methodologies employed in the conjoint experiment literature, as discussed by Hainmueller, Hopkins and Yamamoto (2014). It is worth noting that the formal derivation of the ITE hinges on other assumptions: Stability and No Carryover Effects, and No Profile-Order Effects, which are standard in the conjoint experiment literature. As shown by Abramson et al. (2023), the Marginal Component Effect (MCE) is defined as a weighted sum of ITEs over different profile combinations in experiments: $MCE_{i}=ITE_{i}\times\mathbb{P}[\text{profiles contain }x_{1}\text{and}\tilde{x}_{1}\text{occur}]$. Then, the Average Marginal Component Effect (AMCE) represents the expected value of MCE across the distribution of DMs: $AMCE=\mathbb{E}[MCE_{i}]$. A more formal discussion and derivation are available in the Supplementary Information (SI) 3.
@@ 2 Attention Effect
Individuals have limited attention. Economists have observed that when people make decisions, such as purchasing a computer, they do not consider all possible options (Hausman, 2008). Within the literature, the subset of alternatives that a DM considers when making a decision is referred to as the “consideration set”, the set of attributes that DM indeed considers when making decisions (Hausman, 2008; Wright and Barbour, 1977). It has been well-argued and demonstrated that due to cognitive limitations, humans cannot allocate sufficient attention to every potential attribute (Stigler, 1961; Jones and Baumgartner, 2005; Chetty, Looney and Kroft, 2009). This indicates that the consideration set is typically smaller than the full set $X$ that may influence a decision. We denote the consideration set in real-world decision-making as $X^{r}\subset X$ and the consideration set in an experimental context as $X^{e}\subset X$.
@@
The consideration set utilized in experiments, denoted as $X^{e}$, is often a subset of the real-world consideration set, $X^{r}$, for several reasons. Firstly, researchers seldom possess complete knowledge of which attributes constitute the real-world consideration set $X^{r}$. Secondly, even with such knowledge, it is impractical to present all relevant information about these attributes within the confines of a single survey experiment. A more general assumption could be that $X^{e}\neq X^{r}$. However, the empirical results in the paper provide little evidence on this assumption.
@@ 2.1 Inconsistency of Effect Magnitude
To compare the two causal effects when alternatives have different attributes (and without a salience effect), we need a benchmark assumption that the salience ratio for any given dimension remains constant.
@@
This assumption reflects the natural idea that if, suppose a DM faces single one profile, DM considers attribute $j$ (e.g., gender) to be more salient than attribute $k$ (e.g., age), that is, $\alpha_{j}>\alpha_{k}$, then in another similar circumstance, again when DM faces only one profile, the same DM should still perceive gender as more salient than age, $\alpha^{\prime}_{j}>\alpha^{\prime}_{k}$, even if the exact values of salience could differ ($\alpha_{j}\neq\alpha^{\prime}_{j}$ and $\alpha_{k}\neq\alpha^{\prime}_{k}$). This holds true particularly because the attributes are uncorrelated and the DM evaluates only one profile at a time, eliminating any influence from other profiles. Thus, there is no reason for the order of salience to reverse. It is important to note that this assumption does not pertain to the consistency of relative importance, which is defined based on causal effects, not salience.
@@
We will test this hypothesis in the following subsection. Specifically, following common practice, the experimental causal effect is measured by the AMCE. As noted before, since AMCE is essentially a weighted sum of ITE, the amplification bias of ITE is also reflected in AMCE.
@@ 3 Salience Effect
In this section, we explore the inconsistency beyond the assumption of fixed salience in the previous section. We adopt a psychological framework developed by Bordalo, Gennaioli and Shleifer (2012) to elucidate how variations in attribute salience can influence this relationship. This approach posits that individuals disproportionately weight more salient attributes when evaluating alternatives.